OpenStudy (anonymous):

solve the quadratic equation by completing square x^2 +8x =7 a. x=-1 +/- sqrt23 b. x=4+sqrt23 c. x=-4+/- 2 sqrt23 d. x=-4+/-sqrt23 a. z=-8+/-sqrt17 b. z=8+sqrt17 c. z= -16 +sqrt47 d. z=8 +/- sqrt47

OpenStudy (anonymous):

b=8 so... (8/2)^2 = 4^2 = 16 so add that to both sides so x^2 + 8x + 16 = 7 + 16 so now you have x^2 + 8x + 16 = 23 so FOIL the left side and you get (x + 4)(x + 4) = 23 or (x + 4)^2 = 23 so x + 4 = sq. root 23 so x = -4 + sq. root 23

OpenStudy (anonymous):

if you dont mind me asking how is it you got the sixteen? you multiplied by two i beleive how is it you got that number?

OpenStudy (anonymous):

ok so the coefficient of x is 8 right? your original equation is x^2 + 8x = 7, so the standard form for that equation is ax^2 + bx + c = 0 well in this particular equation there is no c cuz c has already been put over on the right side of the equation, and a=1 since there's nothing in front of x^2...and there's an 8 in front of x so b=8. so you take half of whatever b is and then square that number....so since b=8 you take (8/2) which is 4 and then square it so 4^2 = 16 now add that number (16) to both sides of your original equation so x^2 + 8x + 16 = 7 + 16 so x^2 + 8x + 16 = 23 understand?

OpenStudy (anonymous):

yes thanks a bunch

OpenStudy (anonymous):

bovice?